Q Two particles A and B are at the same point on the circumference of a circle when they start moving along the circumference in opposite directions at 35 degree per second and 25 degree per second respectively. how many times do they cross each other 
  1. in one minute 
  2. in 95 seconds

Grade:11

1 Answers

Vikas TU
7 years ago
Make one of the particle at rest and provide its angular velocity to another one in opposite direction from its own.
Let’s A be at rest 
then the angular vel. of particle B will be =>
60 degree/second clockwise (say)

a) in a minute
there are 60 seconds and hence 360 degree.
Hence crossing each other =  360/60 = 6 times.

b) 95 seconds = 1.58 minute => 568.8 degree.

no. of times crossing each other = > 568.8/60 =  9.48 =>>> 10 times crossing.